OpenStudy (31356):

If the polynomial cannot be factored, write PRIME Factoring Using Distributive Property. MEDALS REWARDED!!

OpenStudy (31356):

x^2-16 Can this be factored?

OpenStudy (31356):

@satellite73 @jim_thompson5910

OpenStudy (anonymous):

Yes. You can use the factorization formula for a difference of two squares. a^2-b^2=(a+b)(a-b) (x+4)(x-4)

OpenStudy (31356):

Okay thanks how about this one?

OpenStudy (31356):

2x^3-128x

OpenStudy (anonymous):

We can factor out a 2 to begin with. 2(x^3-64x) Now it's a difference of two cubes. There's a factorization formula for that, as well: a^3-b^3=(a-b)(a^2+ab+b^2) We can factor 2(x^3-64x) using that. 2(x-4)(x^2+4x+16)